# Student-AI Interaction: A Case Study of CS1 students

### v1

@@ Abstract.
The new capabilities of generative artificial intelligence tools (Generative AI), such as ChatGPT, allow users to interact with the system in intuitive ways, such as simple conversations, and receive (mostly) good-quality answers. These systems can support students’ learning objectives by providing accessible explanations and examples even with vague queries. At the same time, they can encourage undesired help-seeking behaviors by providing solutions to the students’ homework. Therefore, it is important to better understand how students approach such tools and the potential issues such approaches might present for the learners.
@@
Our results suggest that in about a third of the cases, the student attempted to complete the task by submitting the full description of the tasks to ChatGPT without making any effort on their own. We also observed that few students verified their solutions. We discuss the results and their potential implications.
@@ 1. Introduction
Many interactive learning environments (ILEs) offer on-demand help, which refers to information actively solicited by the learner and provided by the system, intended to positively influence learning (Aleven et al., 2003). While effective help-seeking behavior in ILEs is linked to improved learning outcomes, many learners do not effectively utilize available help resources(Aleven and Koedinger, 2000; Luckin et al., 1999). This raises concerns that ILEs may not reach their full potential unless we find ways to help students use these support tools better. Additionally, since seeking help is a crucial skill that affects learning in many situations(Nelson-Le Gall, 1981; Newman, 2023; Aleven and Koedinger, 2001), designing ILEs to encourage effective help-seeking behaviors could greatly enhance their educational value(Aleven and Koedinger, 2001; Gräsel et al., 2000).
@@
Generative AI can be considered an interactive learning environment (ILE), as they can provide on-demand help to students, such as providing explanations of concepts and offering examples. They play a crucial role as interactive components within broader educational settings. They contribute to interactive learning by providing immediate responses and facilitating learning interactions in real time. With the advance of Generative AI tools such as ChatGPT and Copilot, they have become integral tools for many students, particularly those in introductory programming courses(Prather et al., 2023a). They can assist students by providing code suggestions, explaining programming concepts, identifying and resolving errors, and generating code documentation (Kazemitabaar et al., 2023b; Denny et al., 2024c). Their widespread adoption signifies a major shift in how programming is learned and practiced. However, this increasing reliance on Generative AI tools presents significant challenges. One primary concern is that students might become overly dependent on these tools, potentially hindering their ability to develop fundamental programming skills (Prather et al., 2024). Additionally, AI-generated solutions might lead to academic dishonesty or reduced problem-solving capabilities among students (Becker et al., 2023). This dichotomy has sparked a debate among educators: while some argue against the use of Generative AI in learning due to these risks, others advocate for its judicious use to enhance learning outcomes (Lau and Guo, 2023).
@@
Despite various studies exploring the impact of Generative AI tools like GitHub Copilot on productivity (Ziegler et al., 2022a; Kazemitabaar et al., 2023a) and learning outcomes (Tan et al., 2024; Margulieux et al., 2024), there remains a critical gap in understanding how novice programmers interact with these tools without any limitations on using Generative AI, where they can freely utilize AI-driven assistance and guidance. Understanding students’ level of trust in these tools, which impacts their adoption and learning outcomes (Amoozadeh et al., 2024, 2023), and analyzing their interactions while using Generative AI for programming tasks is crucial. These insights can help us identify behaviors, actions, decisions, and responses while utilizing Generative AI tools. By modeling their behaviors and identifying challenges, we can develop systems that provide effective AI-assisted learning, thereby enabling students to learn programming concepts more efficiently and achieve better educational outcomes.
@@
To address this gap, our study investigates the interaction of novice programmers with ChatGPT while solving programming tasks. We conducted a mixed-method study involving 15 CS1 students using a specially designed VSCode plugin that integrates ChatGPT directly into their coding environment. This setup allows us to observe their natural use of Generative AI assistance without external constraints. Our study aims to understand how, when, and why students use ChatGPT, and what strategies they employ to integrate Generative AI into their problem-solving processes.
This paper aims to better understand how students use ChatGPT during programming tasks, revealing diverse usage patterns and their effectiveness. In particular, we seek to answer the following research questions:
@@
    How frequently and at what stages do students use ChatGPT while solving programming tasks?
@@
    How do students interact with ChatGPT while solving a programming task?
@@
    What strategies do novices employ when interacting with ChatGPT?
Our findings highlight that participants, such as P2, extensively interacted with Generative AI yet faced challenges in achieving accurate solutions. Others adopted approaches including step-by-step prompting or hybrid approaches that combine independent coding and Generative AI support. The acceptance of the Generative AI responses varied, from full adoption of the Generative AI’s response to using Generative AI to find answers to the queries, or simply comparing the Generative AI responses to their own solutions. Collaboration success rates showed that Generative AI facilitated correct answers in 65% of the cases, while 34% remained unresolved despite using Generative AI.
2. Related Work
2.1. Exploring the Impact of GenAI in Programming
The recent widespread deployment of generative AI programming assistants has motivated much research on the use of these tools. Researchers have conducted empirical studies (Leinonen et al., 2023; Hellas et al., 2023; Sarsa et al., 2022) to evaluate the quality of code and explanations that GenAI programming assistants can generate. Others have explored the usefulness of LLM-based programming tools through user studies (Ziegler et al., 2022b; Vaithilingam et al., 2022; Barke et al., 2023; kazemitabaar2023; Xu et al., 2022; Mozannar et al., 2022; Ross et al., 2023; Srinivasa Ragavan and Alipour, 2024). Notably, Vaithilingam (Vaithilingam et al., 2022) compared the user experience of GitHub Copilot with traditional auto-complete and found that programmers faced more frequent difficulties in completing tasks with Copilot, although there was no significant impact on task completion time. Barke (Barke et al., 2023) took a step further to understand how programmers interact with code-generating models using a grounded theory analysis, identifying two primary modes of interaction: acceleration, where GenAI is used to speed up code authoring in small logical units, and exploration, where GenAI serves as a planning assistant by suggesting structure or API calls. While these studies have contributed to our understanding of how developers use GenAI developer tools, they mainly focused on evaluating the effectiveness of such tools and did not directly address developers’ perceptions of them. Some studies, such as those by Liang (Liang et al., 2023) and Ziegler (Ziegler et al., 2022b), have also focused on how developers perceive these tools. For instance, Liang (Liang et al., 2023) conducted a survey involving 410 developers to investigate the usability challenges associated with GenAI programming assistants, finding that although developers appreciate their auto-complete capabilities, they often face challenges in controlling the quality of code generation and have concerns about potentially infringing on intellectual property. Although these studies provide valuable insights on how users perceive the GenAI tools, the insights are primarily applicable to professional programmers.
2.2. The Role of AI in Education
Recent studies have started to explore these GenAI tools’ applications in educational contexts. For example, Denny et al. (Denny et al., 2024b) have identified desirable characteristics for GenAI teaching assistants in programming education. They emphasize the importance of features that provide instant, engaging support while allowing students to maintain autonomy in their learning process. Liu et al. (Liu et al., 2024) investigated the integration of GenAI tools in the CS50 course at Harvard University, demonstrating how generative GenAI can significantly enhance the learning experience for introductory programming students. Their approach involved developing and deploying a suite of GenAI-based tools designed to emulate a 1:1 teacher-to-student ratio, thereby providing personalized, real-time support to students. These tools, including the CS50 Duck, a virtual rubber-duck debugging assistant, were well-received by students, who reported that the GenAI tools made them feel as though they had a personal tutor available at all times. Prather et al. (Prather et al., 2023b) have recently published a series of studies examining the usability and interaction challenges faced by novice programmers using GenAI tools like GitHub Copilot. Their research highlights the dual nature of such tools: while they can significantly accelerate the coding process and aid in overcoming programming blocks, they also introduce unique cognitive and metacognitive difficulties. Students often found it ”weird” how accurately Copilot predicted their needs, which led to a mix of fascination and dependency concerns. The study also identified new interaction patterns such as ”shepherding” and ”drifting,” reflecting the nuanced ways novices attempt to guide the GenAI or get misled by it. These insights underscore the importance of designing GenAI tools that support learning while promoting independent problem-solving skills and managing over-reliance.
2.3. Interaction Patterns of Novice Programmers Using LLMs
Kazemitabaar et al. (Kazemitabaar et al., 2023b) conducted a thematic analysis of novice learners using an LLM-based code generator in a self-paced Python course. They identified distinct coding approaches among the learners, noting that the Hybrid approach, which combined manual coding with LLM assistance, produced the best outcomes. However, the study also highlighted signs of over-reliance on LLMs, such as copying LLM output without changes, and positive self-regulation behaviors, like adding code to verify LLM output. Nguyen et al. (Nguyen et al., 2024) conducted a large-scale, multi-institution study that explored the challenges faced by near-novices when interacting with Code LLMs. Their study identified barriers such as difficulties in expressing problem understanding and using appropriate coding terminology. Prather et al. (Prather et al., 2024) further investigated the impact of generative AI tools on novice programmers’ metacognitive awareness and problem-solving strategies, highlighting the potential widening gap between well-prepared and under-prepared students in the era of generative AI.
2.4. Assessing Student Self-Efficacy in AI-Enhanced Learning
Self-efficacy, or the belief in one’s capabilities to achieve a goal or an outcome, is a crucial factor in students’ learning processes. Recent studies have examined how interactions with GenAI tools influence students’ self-efficacy in programming. Tankelevitch et al. (Tankelevitch et al., 2024) discussed the metacognitive demands imposed by GenAI systems, highlighting the need for tools that support students in monitoring and controlling their learning processes. Xue et al. (Xue et al., 2024) investigated the impact of ChatGPT on introductory programming students, finding that the AI tool can enhance students’ self-efficacy by providing immediate feedback and support, thereby reducing the anxiety associated with complex problem-solving tasks. Furthermore, Denny et al. (Denny et al., 2024b) emphasized the importance of designing AI teaching assistants that not only assist with immediate problem-solving but also help students develop long-term self-efficacy by encouraging independent learning and critical thinking.
3. Methodology
The methodology of this study is geared towards examining how CS1 students interact with generative AI tools in a safe environment without limitations on using Generative AI. We specifically chose to implement a ChatGPT plug-in within the VSCode environment. Using ChatGPT within VSCode aims to reduce the cognitive load associated with switching between different environments, thereby allowing participants to remain within a single, cohesive coding environment.
Tankelevitch et al. (Tankelevitch et al., 2024) research highlights the metacognitive demands placed on users when interacting with generative Generative AI systems. These demands include the need for self-awareness of task goals, task decomposition, and evaluating Generative AI-generated outputs (Tankelevitch et al., 2024). This design choice was influenced by studies that suggest minimizing context switching can enhance user focus and reduce metacognitive demands (Tankelevitch et al., 2024). Within this setup, users encounter a question box where they can freely input prompts without any restrictions on the number or type of questions, as shown in Figure 1 ”Ask a question”.
We recruited 15 CS1 students to participate in a user study supervised by one of our authors in an office setting. Each participant was offered bonus points for their involvement. We considered offering bonus points because we did not want to put any pressure on them. We also clarified that their performance would not be judged by others to ensure that the participation was non-coercive. Participants were given the option to utilize Generative AI assistance at their discretion or complete tasks independently.
3.1. Survey Questions
We asked participants to complete a survey before beginning the programming tasks, referred to as the pre-study, and another survey, known as the post-study, after completing the programming tasks. The pre-study and post-study included questions on self-efficacy (Pintrich et al., 1991). Our purpose is to compare participants’ levels of self-efficacy before and after programming, particularly among those who used Generative AI to solve programming tasks. We calculated the mean self-efficacy score for each participant to compare their levels before and after the study. The pre-study was divided into four sections, whereas the post-study solely comprised self-efficacy and demographic questions.
@@ 3.2. Participants
In this study, we enrolled a total of 15 participants, comprising 8 females and 7 males. Among them, two participants were first-generation students. Within this group of 15 students, 13 were freshmen, one was a junior, and one was a post-baccalaureate student. All participants were English speakers as their first or second language, and we did not encounter any reported difficulties in explaining concepts or instructions in English. Table 1 shows the participants’ demographics.
Participant Age Gender Education-level Racial/Ethnic First-generation
P1 19 Female Freshman African American No
P2 18 Female Freshman African American No
P3 19 Female Freshman African American No
P4 19 Male Freshman Asian No
P5 18 Female Freshman Asian No
P6 19 Female Freshman Asian Yes
P7 19 Male Freshman African American No
P8 18 Male Freshman Latino Yes
P9 22 Female Postbacc African American No
P10 18 Male Freshman Asian No
P11 19 Female Freshman White No
P12 19 Male Freshman African American No
P13 18 Male Freshman African American No
P14 20 Male Junior Latino No
P15 19 Female Freshman African American No
@@
3.3. Protocol
To ensure ethical compliance, the informed consent form for this research study was created in adherence to the guidelines set by the Institutional Review Board (IRB) at a large public university in the US. The informed consent form provided participants with essential information about the study, such as its voluntary nature, the estimated time required (here, no more than 60 minutes), and the assurance of confidentiality. It emphasized that no identifiable information would be collected or stored, and participants were given the option to provide their email addresses separately for future study purposes. The form highlighted that there were no expected risks associated with participating in the research and encouraged participants to ask any questions they had before deciding to participate.

### v2

@@ Abstract.
Generative artificial intelligence tools (Generative AI), such as ChatGPT, allow users to interact with them in intuitive ways (e.g., conversational) and receive (mostly) good-quality answers. In education, such systems can support students’ learning objectives by providing accessible explanations and examples even when students pose vague queries. But, they also encourage undesired help-seeking behaviors, such as by providing solutions to the students’ homework. Therefore, it is important to better understand how students approach such tools and the potential issues such approaches might present for the learners.
@@
Our results suggest that, in about a third of the cases, the student attempted to complete the task by submitting the full description of the tasks to ChatGPT without making any effort on their own. We also observed that few students verified their solutions. We discuss the potential implications of these results.
Generative Artificial Intelligence, Human-AI Interaction, Self-regulation, CS1, User study, Novice programmers
ccs: Human-centered computing Human computer interaction (HCI)ccs: Social and professional topics Computing educationccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Many interactive learning environments (ILEs) offer on-demand help to positively influence learning; here, the learner actively seeks information, and the systems provide the same (Aleven et al., 2003). Although such forms of help-seeking behaviors in ILEs, when effective, are linked to improved learning outcomes, many learners do not use the available help resources effectively(Aleven and Koedinger, 2000; Luckin et al., 1999). This raises concerns that ILEs may not reach their full potential unless we find ways to help students use these support tools better. Furthermore, since seeking help is a crucial skill that affects learning in many situations(Nelson-Le Gall, 1981; Newman, 2023; Aleven and Koedinger, 2001), designing ILEs to encourage effective help-seeking behaviors could greatly enhance their educational value(Aleven and Koedinger, 2001; Gräsel et al., 2000).
@@
Generative AI can be considered an interactive learning environment (ILE), as they can provide on-demand help, such as explanations of concepts and working code examples, to students. They play a crucial role as interactive components within broader educational settings, and contribute to interactive learning by providing immediate responses and facilitating learning interactions in real time. Generative AI tools such as ChatGPT and Copilot have become integral tools for many students, particularly those in introductory programming courses (Prather et al., 2023a). They can assist students by providing code suggestions, explaining programming concepts, identifying and resolving errors, and generating code documentation (Kazemitabaar et al., 2023c; Denny et al., 2024c). Their widespread adoption signifies a major shift in how programming is learned and practiced. However, this increasing reliance on Generative AI tools presents significant challenges. One primary concern is that students might become overly dependent on these tools, potentially hindering their ability to develop fundamental programming skills (Prather et al., 2024b). Additionally, AI-generated solutions might lead to academic dishonesty or reduced problem-solving capabilities among students (Becker et al., 2023). This dichotomy has sparked a debate among educators: while some argue against the use of Generative AI in learning due to these risks, others advocate for its judicious use to enhance learning outcomes (Lau and Guo, 2023).
@@
Despite various studies exploring the impact of Generative AI tools like GitHub Copilot on productivity (Ziegler et al., 2022a; Kazemitabaar et al., 2023a) and learning outcomes (Tan et al., 2024; Margulieux et al., 2024), there remains a critical gap in understanding how novice programmers interact with these tools without any limitations on using Generative AI, where they can freely utilize AI-driven assistance and guidance. Understanding the level of trust of students in these tools, which impacts their adoption and learning outcomes (Amoozadeh et al., 2024, 2023), and analyzing their interactions while using Generative AI for programming tasks is crucial. These insights enable us to identify students’ behaviors, actions, decisions, and responses when using Generative AI tools, guiding the design of AI-assisted learning interfaces that enhance conceptual understanding and drive improved learning outcomes.
@@
To this end, our study investigates the interaction of novice programmers with ChatGPT while solving programming tasks. We conducted a mixed-method study involving 15 CS1 students who used a custom VSCode plugin that integrated ChatGPT directly into their coding environments. This setup allowed us to observe their natural use of Generative AI assistance without external constraints, and to gain a deeper understanding of how students use Generative AI in completing programming tasks, namely 1) when, how and why they use them and to what effectiveness, and 2) what strategies they employ to integrate Generative AI into their problem-solving processes.
Our findings highlight that participants extensively interacted with Generative AI, yet successfully provided correct answers only in 65% of the cases; the rest remained unsolved. Some successful interactions involved step-by-step prompting, or hybrid approaches that combine independent programming and Generative AI support. The overall acceptance of the Generative AI responses also varied, from full adoption of the Generative AI’s response to using Generative AI to find answers to the queries, or simply comparing the Generative AI responses to their own solutions.
2. Related Work
2.1. Exploring the Impact of Generative AI in Programming
The recent widespread deployment of generative AI programming assistants has motivated much research on the use of these tools. Researchers have conducted empirical studies (Leinonen et al., 2023; Hellas et al., 2023; Sarsa et al., 2022) to evaluate the quality of code and explanations that Generative AI programming assistants can generate. Others have explored the usefulness of Large Language Model (LLM)-based programming tools through user studies (Ziegler et al., 2022b; Vaithilingam et al., 2022; Barke et al., 2023; Kazemitabaar et al., 2023b; Xu et al., 2022; Mozannar et al., 2022; Ross et al., 2023; Srinivasa Ragavan and Alipour, 2024). Notably, Vaithilingam et al. (Vaithilingam et al., 2022) compared the user experience of GitHub Copilot with traditional auto-complete and found that programmers faced more frequent difficulties in completing tasks with Copilot, although there was no significant impact on task completion time. Barke et al. (Barke et al., 2023) took a step further to understand how programmers interact with code-generating models using a grounded theory analysis, identifying two primary modes of interaction: acceleration, where Generative AI is used to speed up code authoring in small logical units, and exploration, where Generative AI serves as a planning assistant by suggesting structure or API calls. While these studies have contributed to our understanding of how developers use Generative AI developer tools, they mainly focused on evaluating the effectiveness of such tools and did not directly address developers’ perceptions of them. Some studies, such as those by Liang et al. (Liang et al., 2023) and Ziegler et al. (Ziegler et al., 2022b), have also focused on how developers perceive these tools. For instance, Liang et al. (Liang et al., 2023) conducted a survey involving 410 developers to investigate the usability challenges associated with Generative AI programming assistants, finding that developers appreciate their autocomplete capabilities, but also reported challenges ranging from the quality of code generation to potentially infringing on intellectual property. Although these studies provide valuable insights on how users perceive the Generative AI tools, the insights are primarily applicable to professional programmers; our study focuses on the use of such tools by CS students, and more studies are needed in this area.
2.2. The Role of AI in Education
Recent studies have started to explore the applications of Generative AI tools in educational contexts. For example, Denny et al. (Denny et al., 2024b) have identified desirable characteristics for Generative AI teaching assistants in programming education. They emphasize the importance of features that provide immediate, engaging support while allowing students to maintain autonomy in their learning process. Liu et al. (Liu et al., 2024) investigated the integration of Generative AI tools in the CS50 course at Harvard University, demonstrating how they can significantly enhance the learning experience for students in introductory programming classes. Their approach involved developing and deploying a suite of Generative AI-based tools designed to emulate a 1:1 teacher-to-student ratio, thereby providing personalized, real-time support to students. These tools, including the CS50 Duck, a virtual rubber-duck debugging assistant, were well received by students, who reported that the Generative AI tools made them feel as if they had a personal tutor available at all times. Prather et al. (Prather et al., 2023b) have recently published a series of studies that examine the usability and interaction challenges faced by novice programmers using Generative AI tools, e.g., GitHub Copilot. Their research highlights the dual nature of such tools: while they can significantly accelerate the coding process and aid in overcoming programming blocks, they also introduce unique cognitive and metacognitive difficulties. Students often found it ”weird” how accurately Copilot predicted their needs, which led to a mix of fascination and dependency concerns. The study also identified new interaction patterns such as ”shepherding” and ”drifting,” reflecting the nuanced ways novices attempt to guide Generative AI or are misled by it. All these prior works together lend motivation to our research, highlighting the potential usefulness of Generative AI tools, while underscoring the need to balance learning support with independent problem-solving skills and over-reliance on Generative AI.
2.3. Interaction Patterns of Novice Programmers Using LLMs
To develop Generative AI systems that balance fostering independent thinking with learning support and avoiding overreliance, it is necessary to understand how students use LLMs in practice. In the past, Kazemitabaar et al. (Kazemitabaar et al., 2023c) conducted a thematic analysis of novice learners using an LLM-based code generator in a self-paced Python course. They identified distinct coding approaches among the learners, noting that the Hybrid approach, which combined manual coding with LLM assistance, produced the best outcomes. However, the study also highlighted signs of over-reliance on LLMs, such as copying LLM output without changes, and positive self-regulation behaviors, like adding code to verify LLM output. In another study, Nguyen et al. (Nguyen et al., 2024) conducted a large-scale multi-institutional study that explored the challenges faced by near-novices when interacting with Code LLMs. Their study identified barriers such as difficulties in expressing problem understanding and using appropriate coding terminology. Prather et al. (Prather et al., 2024b) further investigated the impact of generative AI tools on novice programmers’ metacognitive awareness and problem-solving strategies, highlighting the potential widening gap between well-prepared and under-prepared students in the era of generative AI. However, there is a lack of studies elucidating when and how students use LLMs, to what effectiveness, and to what effect on their self-efficacy; this is what we cover in the present study.
2.4. Assessing Student Self-Efficacy in AI-Enhanced Learning
Self-efficacy, or the belief in one’s capabilities to achieve a goal or an outcome, is a crucial factor in students’ learning processes. Recent studies have examined how interactions with GenAI tools influence students’ self-efficacy in programming. Tankelevitch et al. (Tankelevitch et al., 2024) discussed the metacognitive demands imposed by GenAI systems, highlighting the need for tools that support students in monitoring and controlling their learning processes. Xue et al. (Xue et al., 2024) investigated the impact of ChatGPT on introductory programming students, finding that the AI tool can enhance students’ self-efficacy by providing immediate feedback and support, thereby reducing the anxiety associated with complex problem-solving tasks. Furthermore, Denny et al. (Denny et al., 2024b) emphasized the importance of designing AI teaching assistants that not only assist with immediate problem-solving but also help students develop long-term self-efficacy by encouraging independent learning and critical thinking. This study serves to add further evidence on the effect of Generative AI use on student self-efficacy, albeit using a different approach (namely, pre- and post-task comparisons).
3. Methodology
This study focuses on students’ use of Generative AI–covering the whats, whys, whens, and hows of doing so. The specific research questions for our study are as follows:
@@ 3. Methodology
    How frequently do students use Generative AI while solving programming tasks?
@@
    How do students interact with Generative AI while solving a programming task?
@@
    How does the self-efficacy of students change before and after programming with Generative AI?
To answer these questions, we adopted a mixed-methods approach, gathering data from 15 CS1 students using Generative AI to perform a programming task. As we reason later in this section, the access to Generative AI in our study was via a plug-in that we designed for the study.
@@ 3.1. Participants
We recruited students from a CS1 course in a large US public university. We sent an email to the course email list, via the instructor (also an author of this paper), inviting volunteers to participate in a research study involving programming tasks with AI. Each participant was offered five bonus points in their course for their participation. We considered offering bonus points so as to not coerce students into participating. Participants could simply fill in an available time slot for the study if they were willing to participate. The course had 110 students, of which 15 students volunteered to participate.
Table 1 summarizes the demographics of the participants. Our study was gender balanced (8 females, 7 males). Two of our participants were first-generation students (i.e., no parent or guardian possessed a four-year college degree); the rest were continuing-generation students and had at least one parent or guardian with a four-year college degree. Most of our participants (13 out of 15) were freshmen; the other two were junior and post-baccalaureate respectively. The participants reported diverse racial/ethnic backgrounds, and everyone used English as their first or second language fluently.
Participant Age Gender Education-level Racial/Ethnic College-generation
P1 19 Female Freshman African American Continuing-generation
P2 18 Female Freshman African American Continuing-generation
P3 19 Female Freshman African American Continuing-generation
P4 19 Male Freshman Asian Continuing-generation
P5 18 Female Freshman Asian Continuing-generation
P6 19 Female Freshman Asian First-generation
P7 19 Male Freshman African American Continuing-generation
P8 18 Male Freshman Latino First-generation
P9 22 Female Postbacc African American Continuing-generation
P10 18 Male Freshman Asian Continuing-generation
P11 19 Female Freshman White Continuing-generation
P12 19 Male Freshman African American Continuing-generation
P13 18 Male Freshman African American Continuing-generation
P14 20 Male Junior Latino Continuing-generation
P15 19 Female Freshman African American Continuing-generation